a swimmers speed in the direction of flow of river is 16km/hr .Against the direction of flow of river, the swimmers speed is 8km/hr .calculate the swimmers speed in still water and the velocity of flow of the river.
Answer by lwsshak3(6473)

a swimmers speed in the direction of flow of river is 16km/hr .Against the direction of flow of river, the swimmers speed is 8km/hr .calculate the swimmers speed in still water and the velocity of flow of the river.
let x=swimmers speed in still water
let c=velocity of flow of the river.
..
x+c=16
x-c=8
add
2x=24
x=12
c=16-x
c=4
swimmers speed in still water=12 km/hr
velocity of flow of the river=4 km/hr